OpenStudy (anonymous):

Ellen has a bag of candy to share with her friends. If she has 2 butterscotch candies, 3 peppermints, and 5 pieces of toffee, what is the probability that the first two people will randomly choose toffees if they take turns choosing? A.9/10 B.2/9 C.1/2 D.1/5

OpenStudy (cwrw238):

probabilty (toffe first choice) = 5/10 .........................2nd .......... = 4/9 required probability = 5/10 * 4/9 = 20 / 90 = 2/9

OpenStudy (anonymous):

B